### v1

# The algebra of hyperinterpolation-class on the sphere

Abstract
This paper considers the so-called concept of hyperinterpolation-class, i.e., the set of all operators derived from the hyperinterpolation operator on the unit sphere. Concretely, we select four different elements in the hyperinterpolation-class, namely filtered hyperinterpolation, Lasso hyperinterpolation, hard thresholding hyperinterpolation and generalized hyperinterpolation introduced by Dai [10], to explore their algebraic properties on the sphere. Based on the idea of a discrete (semi) inner product, we propose the concepts of hyper self-adjoint operator, hyper projection operator and hyper algebra. Next, we prove generalized hyperinterpolation is hyper self-adjoint and commutative with hyperinterpolation. Then we establish corresponding results of the product, sum and difference of hyper projection operators. Last but not least, we present the results of ideals between hard thresholding hyperinterpolation and hyperinterpolation. We also give a preliminary result of involutions, and also introduce the concepts of hyper $C^{\ast}$-algebra and hyper homomorphism.
Keywords: hyperinterpolation-class, projection, algebra, ideals, involutions
AMS subject classifications. 41A10, 41A36, 47L20, 47L80
1 Introduction
In this paper, we explore the algebraic properties of hyperinterpolation-class on the unit sphere $\mathbb{S}^{d}\subset\mathbb{R}^{d+1}$ and extend concepts and results from functional analysis into the realm of hyperinterpolation. Hyperinterpolation, initially introduced by Sloan in 1995 [28], offers a robust and efficient approximation for continuous functions in high-dimensional settings. It relies on high-order quadrature rules to approximate truncated Fourier expansions corresponding to orthonormal polynomial bases over compact subsets or manifolds [28].
Various variants of hyperinterpolation have been developed [2, 3, 4, 10, 21, 23, 24, 25, 31], broadening the applicability of hyperinterpolation across different scenarios. Recent advancements, such as relaxing or bypassing quadrature exactness assumptions via the Marcinkiewicz-Zygmund property [5, 6], have further enhanced the utility of hyperinterpolation. There are also many works on hyperinterpolation [7, 8, 9, 15, 14, 18, 20, 30, 32, 33, 35]. This paper focuses on the algebraic aspects of hyperinterpolation-class on the sphere $\mathbb{S}^{d}$, introducing concepts like hyper operator norms, hyper self-adjoint operators, and hyper projection operators.
Projection plays a major role in Hilbert spaces for certain problems, such as spectral theorem [27]. Notably, hyperinterpolation operators are identified as projection operators, serving as the unique solution to a weighted least squares approximation problem [28]. Many properties of hyper projection operators are established with the aid of a semi (inner) product, like the product, sum, and difference operations. We also propose the concept of hyper algebra over the set of continuous functions on the sphere, highlighting elements that form an algebra and satisfy the Pythagorean theorem. We prove that hard thresholding hyperinterpolation operators and hyperinterpolation operators belong to the hyper algebra.
Additionally, we delve into ideals and involutions of hyperinterpolation-class, identifying the hyper algebra formed by hard thresholding hyperinterpolation operators as the maximal ideal of a broader hyper algebra consisting of hard thresholding hyperinterpolation and hyperinterpolation operators. We introduce the notion of hyper $C^{\ast}$-algebras based on involutions and demonstrate that the hyper algebra comprising hyperinterpolation operators qualifies as a hyper $C^{\ast}$-algebra. Furthermore, we present the concept of hyper homomorphism and establishe that hyperinterpolation operators serve as hyper homomorphisms within the hyper algebra.
In the next section, we review several fundamental terminologies in spherical approximation and introduce some variants of hyperinterpolation. After mathematical preliminaries in Section 2, a series of definitions pertinent to hyper projection is proposed in Section 3. The ideals and involutions of hyper algebras are defined and analyzed in Section 4.
2 Backgrounds on hyperinterpolation
In this section, we introduce some necessary knowledge in spherical approximation [12, 19, 26].
Let $\mathbb{S}^{d}:=\left\{\mathbf{x}\in\mathbb{R}^{d+1}:\|\mathbf{x}\|_{2}=1\right\}$ with $d\geq 2$ be the unit sphere in $\mathbb{R}^{d+1}$, and let the space $L^{2}=L^{2}(\mathbb{S}^{d})$ be the usual Hilbert space of square-integrable functions on $\mathbb{S}^{d}$ with the inner product
$\langle f,g\rangle_{L^{2}}=\int_{\mathbb{S}^{d}}f(\mathbf{x})g(\mathbf{x}){{{\rm{d}}\omega_{d}(\mathbf{x})}},\quad\forall f,g\in L^{2}(\mathbb{S}^{d}),$
and the induced norm $\|f\|_{L^{2}}:=\langle f,f\rangle^{1/2}_{L^{2}}$, where $\omega_{d}(\mathbf{x})$ denotes the Lebesgue surface measure on $\mathbb{S}^{d}$. The surface area of $\mathbb{S}^{d}$ is denoted by $\omega_{d}$,
$\omega_{d}:=\frac{2\pi^{\frac{d+1}{2}}}{\Gamma\left(\frac{d+1}{2}\right)},$
where $\Gamma(n):=(n-1)!$. Let $\mathbb{P}_{n}(\mathbb{S}^{d})\subset L^{2}(\mathbb{S}^{d})$ be a linear space of all spherical polynomial on $\mathbb{S}^{d}$ of total-degree not exceeding $n$. Note that the restriction of any real harmonic homogeneous polynomial on $\mathbb{R}^{d+1}$ of exact degree $\ell$ to $\mathbb{S}^{d}$ is called a real spherical harmonic of degree $\ell$. Let $\mathbb{H}_{\ell}(\mathbb{S}^{d})$ be the space of spherical harmonics of degree $\ell$ with the dimension $Z(d,\ell):=\dim(\mathbb{H}_{\ell}(\mathbb{S}^{d}))$, given by
$Z(d,0):=1,\quad Z(d,\ell):=\frac{(2\ell+d-1)(\ell+d-2)!}{(d-1)!\ell!},\quad\ell\in\mathbb{N}.$
We denote by
$\left\{Y_{\ell k}^{(d)}:k=1,\cdots,Z(d,\ell)\right\}$ (2.1)
a given $L^{2}(\mathbb{S}^{d})$-orthonormal system of real spherical harmonics of degree $\ell$. Then (2.1) is an $L^{2}(\mathbb{S}^{d})$-orthonormal basis for $\mathbb{H}_{\ell}(\mathbb{S}^{d})$, and the following holds
$\mathbb{P}_{n}(\mathbb{S}^{d})=\mathbb{H}_{0}(\mathbb{S}^{d})\bigoplus\mathbb{H}_{1}(\mathbb{S}^{d})\bigoplus\cdots\bigoplus\mathbb{H}_{n}(\mathbb{S}^{d})$ (2.2)
and
$d_{n}:=\dim(\mathbb{P}_{n}(\mathbb{S}^{d}))=\sum_{\ell=0}^{n}Z(d,\ell)=Z(d+1,n)=\frac{(2n+d)(n+d-1)!}{d!n!}.$
The reproducing kernel $G_{\ell}:\mathbb{S}^{d}\times\mathbb{S}^{d}\to\mathbb{R}$ of the space $\mathbb{H}_{\ell}(\mathbb{S}^{d})$ is
$G_{\ell}(\mathbf{x},\mathbf{y}):=\sum_{k=1}^{Z(d,\ell)}Y_{\ell k}^{(d)}(\mathbf{x})Y_{\ell k}^{(d)}(\mathbf{y})=\frac{2\ell+d-1}{(d-1)\omega_{d}}C_{\ell}^{\frac{d-1}{2}}(\mathbf{x}\cdot\mathbf{y}),$ (2.3)
where $C_{\ell}^{\frac{d-1}{2}}$ is the Gegenbauer polynomial of degree $\ell$ and of index $\frac{d-1}{2}$, and $\mathbf{x}\cdot\mathbf{y}$ is the Euclidean inner product. Then $G_{\ell}$ satisfies the following properties:
$G_{\ell}(\mathbf{x},\cdot)\in\mathbb{P}_{\ell}(\mathbb{S}^{d}),\quad\mathbf{x}\in\mathbb{S}^{d};\qquad G_{\ell}(\mathbf{x},\mathbf{y})=G_{\ell}(\mathbf{y},\mathbf{x}),\quad\mathbf{x},\mathbf{y}\in\mathbb{S}^{d};$
$p(\mathbf{x})=\langle p,G_{\ell}(\mathbf{x},\cdot)\rangle_{L^{2}},\quad\mathbf{x}\in\mathbb{S}^{d},\quad p\in\mathbb{H}_{\ell}(\mathbb{S}^{d}).$
Therefore, by (2.2), we have
$p=\langle p,E_{n}(\mathbf{x},\cdot)\rangle_{L^{2}},\quad\forall\mathbf{x}\in\mathbb{S}^{d},\quad\forall p\in\mathbb{P}_{n}(\mathbb{S}^{d}),$
where
$E_{n}(\mathbf{x},\mathbf{y}):=\sum_{\ell=0}^{n}G_{\ell}(\mathbf{x},\mathbf{y})=\frac{1}{\omega_{d}}\frac{(d)_{n}}{\left(\frac{d}{2}\right)_{n}}P_{n}^{(\frac{d}{2},\frac{d-2}{2})}(\mathbf{x}\cdot\mathbf{y})$ (2.4)
with the Pochhammer symbol (falling factorial)
$(a)_{0}:=1,\quad(a)_{\ell}:=a(a+1)\cdots(a+\ell-1),\quad a\in\mathbb{R},\ell\in\mathbb{N},$
and $P_{n}^{(\frac{d}{2},\frac{d-2}{2})}$ the Jacobi polynomial [34] of degree $n$ with indices $\frac{d}{2}$ and $\frac{d-2}{2}$.
From $L^{2}(\mathbb{S}^{d})$ to the polynomial space $\mathbb{P}_{n}(\mathbb{S}^{d})$, there exists a unique projection $\mathcal{P}_{n}:L^{2}(\mathbb{S}^{d})\to\mathbb{P}_{n}(\mathbb{S}^{d})$, i.e.,
$\mathcal{P}_{n}f:=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\hat{f}_{\ell k}^{(d)}Y_{\ell k}^{(d)}=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{L^{2}}Y_{\ell k}^{(d)},\quad\forall f\in L^{2}(\mathbb{S}^{d}),$
with the Fourier coefficients
$\hat{f}_{\ell k}^{(d)}:=\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{L^{2}}=\int_{\mathbb{S}^{d}}f(\mathbf{x})Y_{\ell k}^{(d)}(\mathbf{x}){{\rm{d}}\omega_{d}(\mathbf{x})}.$ (2.5)
With the help of $E_{n}$ of $\mathbb{P}_{n}(\mathbb{S}^{d})$, we can rewrite the $L^{2}(\mathbb{S}^{d})$-orthogonal projection $\mathcal{P}_{n}$ onto $\mathbb{P}_{n}(\mathbb{S}^{d})$ as
$\mathcal{P}_{n}f(\mathbf{x})=\int_{\mathbb{S}^{d}}f(\mathbf{y})E_{n}(\mathbf{x},\mathbf{y}){\rm{d}}\omega_{d}(\mathbf{y})=\left\langle f,E_{n}(\mathbf{x},\cdot)\right\rangle_{L^{2}},\quad\forall\mathbf{x}\in\mathbb{S}^{d}.$
In the general case, the computation of the integral in (2.5) is a challenging task [13]. Thus, to approximate the Fourier-like coefficients, we seek the help of a quadrature rule
$\sum_{j=1}^{N}w_{j}f(\mathbf{x}_{j})\approx\int_{\mathbb{S}^{d}}f(\mathbf{x}){\rm{d}}\omega_{d}(\mathbf{x}),\quad\forall f\in\mathcal{C}(\mathbb{S}^{d}),$ (2.6)
where $\mathbf{w}=\{w_{1},\ldots,w_{N}\}$ is a set of positive weights and $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\subset\mathbb{S}^{d}$ is a set of nodes, and $\mathcal{C}(\mathbb{S}^{d})$ is the set of all continuous functions over $\mathbb{S}^{d}$. A quadrature rule (2.6) is said to have quadrature exactness $m$ if
$\sum_{j=1}^{N}w_{j}p(\mathbf{x}_{j})=\int_{\mathbb{S}^{d}}p(\mathbf{x}){\rm{d}}\omega_{d}(\mathbf{x}),\quad\forall p\in\mathbb{P}_{m}(\mathbb{S}^{d}).$
Under a quadrature rule (2.6) with quadrature exactness $2n$, we introduce the “discrete (semi) inner product[28]
$\langle f,g\rangle_{N}:=\sum_{j=1}^{N}w_{j}f(\mathbf{x}_{j})g(\mathbf{x}_{j}),$ (2.7)
for any $f,g\in\mathcal{C}(\mathbb{S}^{d})$. Then we refer to the seminal paper [28], where Sloan introduced the hyperinterpolation operator $\mathcal{L}_{n}:\mathcal{C}(\mathbb{S}^{d})\to\mathbb{P}_{n}(\mathbb{S}^{d})$ as
$\mathcal{L}_{n}f:=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\left\langle{f,Y_{\ell k}^{(d)}}\right\rangle_{N}Y_{\ell k}^{(d)},\qquad\forall f\in\mathcal{C}(\mathbb{S}^{d}).$ (2.8)
With the aid of the reproducing kernel (2.4), for $f\in\mathcal{C}(\mathbb{S}^{d})$, we can express the hyperinterpolant $\mathcal{L}_{n}f\in\mathbb{P}_{n}(\mathbb{S}^{d})$ as the following
$\mathcal{L}_{n}f(\mathbf{x})=\langle f,E_{n}(\mathbf{x},\cdot)\rangle_{N}=\sum_{j=1}^{N}w_{j}f(\mathbf{x}_{j})E_{n}(\mathbf{x},\mathbf{x}_{j}),\quad\forall\mathbf{x}\in\mathbb{S}^{d}.$
For more properties on hyperinterpolation on the sphere, we refer to [6, 19, 20, 31, 30].
2.1 Variants of hyperinterpolation
To give examples of various algebraic properties of hyperinterpolation, we need to review some variants of hyperinterpolation, named filtered hyperinterpolation [31], Lasso hyperinterpolation [4], hard thresholding hyperinterpolation [2] and generalized hyperinterpolation [10]. Differing from those single operators variants of hyperinterpolation, we constructed a variant named hybrid hyperinterpolation which is a composition of filtered hyperinterpolation and Lasso hyperinterpolation [3]. In this paper, we mainly investigate single operator varinats of hyperinterpolation.
Now, we first list some important definitions which are crucial in variants of hyperinterpolation. The following definitions of soft and hard thresholding operators are closely related to Lasso hyperinterpolation and hard thresholding hyperinterpolation, respectively.
Definition 2.1 ([16])
The soft thresholding operator, denoted by $\eta_{S}(a,k)$, is defined as
$\eta_{S}(a,k):=\max(0,a-k)+\min(0,a+k).$
Definition 2.2 ([16])
The hard thresholding operator, denoted by $\eta_{H}(a,k)$, is defined as
$\eta_{H}(a,k):=\left\{\begin{array}[]{ll}a,&\text{if}~{}~{}|a|>k,\\ 0,&\text{if}~{}~{}|a|\leq k.\end{array}\right.$
The following is the definition of Lasso hyperinterpolation.
Definition 2.3 ([4])
Given a quadrature rule (2.6) with exactness $2n$, a Lasso hyperinterpolation of $f\in\mathcal{C}(\mathbb{S}^{d})$ onto $\mathbb{P}_{n}(\mathbb{S}^{d})$ is defined as
$\mathcal{L}_{n}^{\lambda}f:=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\eta_{S}\left(\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N},\lambda\mu_{\ell k}\right)Y_{\ell k}^{(d)},$ (2.9)
where $\lambda>0$ is the regularization parameter and $\mu_{\ell k}$’s are a set of positive penalty parameters.
Similarly, we can define hard thresholding hyperinterpolation with the aid of a hard thresholding operator $\eta_{H}(\cdot,\lambda)$.
Definition 2.4 ([2])
Given a quadrature rule (2.6) with exactness $2n$, a hard thresholding hyperinterpolation of $f\in\mathcal{C}(\mathbb{S}^{d})$ onto $\mathbb{P}_{n}(\mathbb{S}^{d})$ is defined as
$\mathcal{H}_{n}^{\lambda}f:=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\eta_{H}\left(\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N},\lambda\right)Y_{\ell k}^{(d)},$ (2.10)
where $\lambda>0$ is the regularization parameter.
Figure 1: Left: the filter function; Right: soft and hard thresholding operators.
Now, we introduce the filter function which occurs in the filtered hyperinterpolation [31].
Definition 2.5
The filter function, denoted by $h(x)$, is defined as
$h(x):=\left\{\begin{array}[]{ll}1,&\text{if}~{}~{}x\in[0,s],\\ 0,&\text{if}~{}~{}x\in[1,\infty),\end{array}\right.$
where $10<s<$ 1 is a constant.
The filter function $h(x)$ has many different forms [29]. In the following, we give two filter functions defined by sin function as examples [1], i.e.,
$h_{1}(x)=\left\{\begin{array}[]{ll}1,&x\in[0,1/2],\\ \sin^{2}(\pi x),&x\in[1/2,1],\\ 0,&x\in[1,+\infty),\end{array}\right.$ (2.11)
and
$h_{2}(x)=\left\{\begin{array}[]{ll}1,&x\in[0,3/4],\\ \sin^{2}(2\pi x),&x\in[3/4,1],\\ 0,&x\in[1,+\infty),\end{array}\right.$ (2.12)
and we display $h_{1}(x)$ and $h_{2}(x)$ on the left of Figure 1.
[…]
Definition 2.6 ([31])
Let ${\overline{{\rm{supp}}(h)}}=[0,a]$, with $0<s<a\leq 1$. Given a quadrature rule (2.6) with exactness $\lceil an\rceil-1+\lfloor n/2\rfloor$, the filtered hyperinterpolant $\mathcal{F}_{n,N}{f}\in\mathbb{P}_{\lceil an\rceil-1}(\mathbb{S}^{d})$ of $f\in\mathcal{C}(\mathbb{S}^{d})$ is defined as
$\mathcal{F}_{n,N}{f}:=\sum_{\ell=0}^{n}h\left(\frac{\ell}{n}\right)\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}.$ (2.13)
where $d_{n}=\dim({\mathbb{P}}_{n}((\mathbb{S}^{d})))$ and $h$ is a filter function defined by Definition 2.5.
For convenience, we use $h_{\ell}$ to represent $h(\ell/n)$ for all $\ell=0,\cdots,n$.
Remark 2.1
With different values $10<s<$ 1 in the filter function $h(x)$, we have
$\mathcal{F}_{n,N}p=p,\quad\forall p\in\mathbb{P}_{\lfloor sn\rfloor}(\mathbb{S}^{d}).$
In the following, we introduce the generalized hyperinterpolation:
Definition 2.7
[10] Let $G_{\ell}(\cdot)$ be defined by (2.3) for $\ell=0,1,\ldots,n$. Assume that
$D_{n}(t)=\sum_{\ell=0}^{n}a_{n\ell}G_{\ell}(t),\quad n=0,1,\ldots,\quad t\in[-1,1],$
is a sequence of polynomials satisfying
$\lim_{n\to\infty}a_{n0}=\lim_{n\to\infty}a_{n1}=1.$
Then the generalized hyperinterpolation operators associated to $\{D_{n}(t)\}_{n=0}^{\infty}$ are defined as
$\mathfrak{G}\mathfrak{L}_{n}f(\mathbf{x})\equiv\mathfrak{G}\mathfrak{L}_{n,D_{n}}f(\mathbf{x})=\sum_{j=1}^{N}w_{j}f(\mathbf{x}_{j})D_{n}(\mathbf{x}\cdot\mathbf{x}_{j})=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}).$ (2.14)
Remark 2.2
Different from the hyperinterpolation operator $\mathcal{L}_{n}$ which demands the quadrature exactness of $2n$, the generalized hyperinterpolation operator $\mathfrak{G}\mathfrak{L}_{n}$ requires that the positive quadrature formula is exact of degree $n+1$ in [10, 25]. In fact, the quadrature exactness $n+1$ is the lowest requirement for generalized hyperinterpolation to guarantee the uniform convergence. The quadrature exactness $2n$ does not influence the uniform convergence for generalized hyperinterpolation [26, Definition 6.9, Theorem 6.34].
Remark 2.3
The generalized hyperinterpolation defined in [25, 26] assumes the positivity of the kernels $D_{n}$. Under this assumption, Reimer proved that the generalized hyperinterpolation operator possesses the important property that the approximation error can be uniformly controlled by the modulus of continuity of the first order [25]. Dai proved the uniform approximation error for $\mathfrak{G}\mathfrak{L}_{n}$ can be controlled by the modulus of continuity of the second order in [10], where he bypassed the assumption of the positivity of the kernels $D_{n}$.
Now, we can define the important concept hyperinterpolation-class:
Definition 2.8
Given a quadrature rule (2.6) with exactness $2n$, for $f\in\mathcal{C}(\mathbb{S}^{d})$, the hyperinterpolation-class on the sphere $\mathbb{S}^{d}$ is a set of operators defined by
$\mathbf{HC}(\mathbb{S}^{d}):=\left\{\mathcal{T}_{n}:\mathcal{T}_{n}f:=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\Phi\left(\left\langle{f,Y_{\ell k}^{(d)}}\right\rangle_{N}\right)Y_{\ell k}^{(d)}\text{ with }\Phi:\mathbb{R}\to\mathbb{R}\text{ being an operator}\right\}.$ (2.15)
Before leaving this subsection we remark that there are many choices of the operator $\Phi:\mathbb{R}\to\mathbb{R}$, such as soft and hard thresholding operators [16]. Lasso, hard thresholding, filtered and generalized hyperinterpolations are elements of the hyperinterpolation-class $\mathbf{HC}(\mathbb{S}^{d})$.
3 Projection
In the field of Functional Analysis, the concept of projection plays a vital role in bounded operators on a Hilbert space [27]. Thus, we give our first interest on the projection properties under the setting of hyperinterpolation. Based on the discrete (semi) inner product (2.7), we introduce the important semi-norm:
Definition 3.1
A real-valued function $\|f\|_{\ell^{2}(\mathbf{w})}$ defined on $\mathbb{S}^{d}$ as the following
$\|f\|_{\ell^{2}(\mathbf{w})}:=\langle f,f\rangle_{N}^{1/2}=\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})|^{2}\right)^{1/2},$ (3.1)
is a semi-norm on $\mathbb{S}^{d}$, where $f\in\mathcal{C}(\mathbb{S}^{d})$, $\mathbf{w}=\{w_{1},\ldots,w_{N}\}$ is a set of positive weights and $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\subset\mathbb{S}^{d}$ is a set of nodes.
Remark 3.1
Note that for a continuous function $f$ on $\mathbb{S}^{d}$, $\|f\|_{\ell^{2}(\mathbf{w})}=0$ does not mean that $f\equiv 0$ because there is a function $f$ that does not vanish in $\mathbb{S}^{d}$ but its values at all nodes $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ are zeros.
Remark 3.2
For the positive homogeneity of $\|\cdot\|_{\ell^{2}(\mathbf{w})}$, we have
$\|\alpha f\|_{\ell^{2}(\mathbf{w})}=\langle\alpha f,\alpha f\rangle_{N}^{1/2}=\left(\sum_{j=1}^{N}\alpha^{2}w_{j}|f(\mathbf{x}_{j})|^{2}\right)^{1/2}=|\alpha|\|f\|_{\ell^{2}(\mathbf{w})},\quad\forall\alpha\in\mathbb{R}.$
Remark 3.3
With the aid of Cauchy-Schwarz inequality, we can verify the triangle inequality of $\|\cdot\|_{\ell^{2}(\mathbf{w})}$. Indeed, for $f,g\in\mathcal{C}(\mathbb{S}^{d})$, we have
$\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|^{2}\leq$ $\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|\cdot|f(\mathbf{x}_{j})|+\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|\cdot|g(\mathbf{x}_{j})|$
$\leq$ $\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|^{2}\right)^{1/2}\cdot\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})|^{2}\right)^{1/2}$
$+\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|^{2}\right)^{1/2}\cdot\left(\sum_{j=1}^{N}w_{j}|g(\mathbf{x}_{j})|^{2}\right)^{1/2},$
which implies that
$\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})+g(\mathbf{x}_{j})|^{2}\right)^{1/2}\leq\left(\sum_{j=1}^{N}w_{j}|f(\mathbf{x}_{j})|^{2}\right)^{1/2}+\left(\sum_{j=1}^{N}w_{j}|g(\mathbf{x}_{j})|^{2}\right)^{1/2}.$
3.1 Hyper projection operators
Sloan revealed that $\mathcal{L}_{n}f$ is the best discrete least squares approximation (weighted by quadrature weights) of $f$ at the quadrature points in [28], which implies the important result as the following:
Lemma 3.1
For a continuous function $f\in\mathcal{C}(\mathbb{S}^{d})$ with hyperinterpolation coefficients $a_{\ell k}=\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N}$ for $\ell=0,\ldots,n$ and $k=1,\ldots,Z(d,\ell)$. Then
$\|f-\mathcal{L}_{n}f\|_{\ell^{2}(\mathbf{w})}\leq\left\|f-\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}c_{\ell k}Y_{\ell k}^{(d)}\right\|_{\ell^{2}(\mathbf{w})}$ (3.2)
for any real numbers $c_{\ell k}$. Moreover, equality holds precisely when $c_{\ell k}=a_{\ell k}$ for all $\ell=0,\ldots,n$ and $k=1,\ldots,Z(d,\ell)$.
Proof.  By direct computation, we have
$f-\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}c_{\ell k}Y_{\ell k}^{(d)}=f-\mathcal{L}_{n}f+\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}b_{\ell k}Y_{\ell k}^{(d)},$
where $b_{\ell k}=a_{\ell k}-c_{\ell k}$. Then by Pythagorean theorem, we complete the proof. $\square$
Before giving some important results, we need to introduce several concepts related to hyper projection defined by Definition 3.5.
Definition 3.2
Let $\mathcal{T}:\mathcal{C}(\mathbb{S}^{d})\to\mathbb{P}_{n}(\mathbb{S}^{d})$ be an operator. We define the hyper operator norm of $\mathcal{T}$ as the following
$\|\mathcal{T}\|_{\ell^{2}(\mathbf{w})}:=\sup_{\|f\|_{\ell^{2}(\mathbf{w})}>0}\frac{\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}}{\|f\|_{\ell^{2}(\mathbf{w})}}=\sup_{\|f\|_{\ell^{2}(\mathbf{w})}=1}\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}.$ (3.3)
Definition 3.3
An operator $\mathcal{T}^{\ast}$ is a hyper adjoint operator of another operator $\mathcal{T}$ in the sense of the discrete (semi) inner product $\langle\cdot,\cdot\rangle_{N}$, i.e.,
$\langle\mathcal{T}f,g\rangle_{N}=\langle f,\mathcal{T}^{\ast}g\rangle_{N},$ (3.4)
for any $f,g\in\mathcal{C}(\mathbb{S}^{d})$.
We mimic self-adjointness in Hilbert spaces [27] for introducing hyper self-adjointness as follows:
Definition 3.4
An operator $\mathcal{T}$ is a hyper self-adjoint operator in the sense of the discrete (semi) inner product $\langle\cdot,\cdot\rangle_{N}$, i.e.,
$\langle\mathcal{T}f,g\rangle_{N}=\langle g,\mathcal{T}f\rangle_{N},$ (3.5)
for any $f,g\in\mathcal{C}(\mathbb{S}^{d})$.
Proposition 3.1
Filtered hyperinterpolation operator $\mathcal{F}_{n,N}$ and hyperinterpolation operator $\mathcal{L}_{n}$ are hyper self-adjoint.
Proof.  For any $f,g\in\mathcal{C}(\mathbb{S}^{d})$, we obtain
$\left\langle\mathcal{F}_{n,N}f,g\right\rangle_{N}$ $=\left\langle\sum_{\ell=0}^{n}h_{\ell}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)},g\right\rangle_{N}$ (3.6)
$=\sum_{\ell=0}^{n}h_{\ell}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y_{\ell k}^{(d)}\right\rangle_{N}\left\langle Y_{\ell k}^{(d)},g\right\rangle_{N}$
$=\left\langle f,\sum_{\ell=0}^{n}h_{\ell}\sum_{k=1}^{Z(d,\ell)}\left\langle Y_{\ell k}^{(d)},g\right\rangle_{N}Y_{\ell k}^{(d)}\right\rangle_{N}=\left\langle f,\mathcal{F}_{n,N}g\right\rangle_{N}.$
If we take $h_{\ell}=1$ for $\ell=0,\ldots,n$ in equation (3.6), then we can obtain
$\left\langle\mathcal{L}_{n}f,g\right\rangle_{N}=\left\langle f,\mathcal{L}_{n}g\right\rangle_{N}.$
Thus, we have completed the proof. $\square$
The following theorem states that the generalized hyperinterpolation operators $\mathfrak{G}\mathfrak{L}_{n}$ are hyper self-adjoint and commutative with the hyperinterpolation operator $\mathcal{L}_{n}$.
Theorem 3.1
Let $f,g\in\mathcal{C}(\mathbb{S}^{d})$. The generalized hyperinterpolation operators $\mathfrak{G}\mathfrak{L}_{n}$ are hyper self-adjoint, i.e.,
$\langle\mathfrak{G}\mathfrak{L}_{n}f,g\rangle_{N}=\langle f,\mathfrak{G}\mathfrak{L}_{n}g\rangle_{N}$
and commutative with $\mathcal{L}_{n}$,
$\mathfrak{G}\mathfrak{L}_{n}(\mathcal{L}_{n}f)=\mathcal{L}_{n}(\mathfrak{G}\mathfrak{L}_{n}f)=\mathfrak{G}\mathfrak{L}_{n}f.$
Proof.  By direct computation, we have
$\langle\mathfrak{G}\mathfrak{L}_{n}f,g\rangle_{N}=\left\langle\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)},g\right\rangle_{N}=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}\langle Y_{\ell k}^{(d)},g\rangle_{N}$
and
$\langle f,\mathfrak{G}\mathfrak{L}_{n}g\rangle_{N}=\left\langle f,\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle g,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)}\right\rangle_{N}=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}\langle g,Y_{\ell k}^{(d)}\rangle_{N}.$
Next, we obtain
$\mathfrak{G}\mathfrak{L}_{n}(\mathcal{L}_{n}f)$ $=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\left\langle\sum_{\ell^{\prime}=0}^{n}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\sum_{\ell^{\prime}=0}^{n}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}\left\langle Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
and
$\mathcal{L}_{n}(\mathfrak{G}\mathfrak{L}_{n}f)$ $=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\left\langle\sum_{\ell^{\prime}=0}^{n}a_{n\ell^{\prime}}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\sum_{\ell^{\prime}=0}^{n}a_{n\ell^{\prime}}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}\left\langle Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}).$
Thus, we have completed the proof. $\square$
Then we can define the hyper projection operator which is parallel with the definition of projection [27].
Definition 3.5
An operator $\mathcal{T}$ is a hyper projection operator, if it satisfies
  1. (i)
    $\mathcal{T}$ is hyper self-adjoint;
  2. (ii)
    $\mathcal{T}^{2}=\mathcal{T}$;
Theorem 3.2
Hyperinterpolation operator $\mathcal{L}_{n}$ is a hyper projection operator.
Proof.  From Proposition 3.1, we know that $\mathcal{L}_{n}$ is hyper self-adjoint. It is easy to check that $\mathcal{L}_{n}(\mathcal{L}_{n})=\mathcal{L}_{n}$. Thus, we complete the proof. $\square$
Remark 3.4
Filtered hyperinterpolation operator $\mathcal{F}_{n,N}$ fails to be a hyper projection operator because $\mathcal{F}_{n,N}\mathcal{F}_{n,N}\neq\mathcal{F}_{n,N}$. Hard thresholding hyperinterpolation operator $\mathcal{H}_{n}^{\lambda}$ is not a hyper projection operator since it is not hyper self-adjoint. Lasso hyperinterpolation operator $\mathcal{L}_{n}^{\lambda}$ is not hyper self-adjoint and idempotent. Thus $\mathcal{L}_{n}^{\lambda}$ is not a hyper projection operator.
Remark 3.5
The generalized hyperinterpolation operators $\mathfrak{G}\mathfrak{L}_{n}$ are not idempotent. Indeed, we have
$\mathfrak{G}\mathfrak{L}_{n}(\mathfrak{G}\mathfrak{L}_{n}f(\mathbf{x}))$ $=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\left\langle\sum_{\ell^{\prime}=0}^{n}a_{n\ell^{\prime}}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}Y_{\ell^{\prime}k^{\prime}}^{(d)}(\mathbf{x}),Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\sum_{\ell^{\prime}=0}^{n}a_{n\ell^{\prime}}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\rangle_{N}\left\langle Y_{\ell^{\prime}k^{\prime}}^{(d)}(\mathbf{x}),Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
$=\sum_{\ell=0}^{n}a_{n\ell}^{2}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)}(\mathbf{x}),$
where $a_{n\ell}=1$ do not hold for all $n=0,1,\ldots,$ and $\ell=0,\ldots,n$. Thus, $\mathfrak{G}\mathfrak{L}_{n}$ can not be a hyper projection operator.
Definition 3.6
For any $f,g\in\mathcal{C}(\mathbb{S}^{d})$, $f$ and $g$ are called hyper orthogonal if
$\langle f,g\rangle_{N}=0.$ (3.7)
In a normed space, the unique decomposition of an element holds [27]. However, the situation is different when we consider the semi-norm $\|\cdot\|_{\ell^{2}(\mathbf{w})}$.
Theorem 3.3
Let the continuous functions space $\mathcal{C}(\mathbb{S}^{d})$ equipped with the semi-norm $\|\cdot\|_{\ell^{2}(\mathbf{w})}$. For any closed subspace $M\subset\mathcal{C}(\mathbb{S}^{d})$, there holds
$M\cap M^{\perp}=\{f:\|f\|_{\ell^{2}(\mathbf{w})}=0,f\in M,f\in M^{\perp}\},$
where $M^{\perp}=\{g:\langle f,g\rangle_{N}=0,f\in M,g\in M^{\perp}\}$.
Proof.  Let $f\in M$ and $f\in M^{\perp}$. Since $\|\cdot\|_{\ell^{2}(\mathbf{w})}$ is a semi-norm, then $\|f\|_{\ell^{2}(\mathbf{w})}^{2}=\langle f,f\rangle_{N}=0$ dose not mean $f\equiv 0$ by Remark 3.1. $\square$
3.2 Hyper algebra
Next, we introduce the hyper algebra $\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ as the following:
Definition 3.7
Let $\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ be a set of all operators $\mathcal{T}:\mathcal{C}(\mathbb{S}^{d})\to\mathbb{P}_{n}(\mathbb{S}^{d})$. We call $\mathfrak{B}(\mathcal{C}(\Omega))$ a hyper algebra if it satisfies the following properties:
  1. (i)
    all elements of $\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ hold
    $\left\{\begin{array}[]{c}\mathcal{T}_{1}(\mathcal{T}_{2}\mathcal{T}_{3})=(\mathcal{T}_{1}\mathcal{T}_{2})\mathcal{T}_{3},\\ (\mathcal{T}_{1}+\mathcal{T}_{2})\mathcal{T}_{3}=\mathcal{T}_{1}\mathcal{T}_{3}+\mathcal{T}_{2}\mathcal{T}_{3},\quad\mathcal{T}_{1}(\mathcal{T}_{2}+\mathcal{T}_{3})=\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{1}\mathcal{T}_{3},\\ \alpha(\mathcal{T}_{1}\mathcal{T}_{2})=(\alpha\mathcal{T}_{1})\mathcal{T}_{2}=\mathcal{T}_{1}(\alpha\mathcal{T}_{2})\end{array}\right.$
    for all $\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ and for all scalars $\alpha$.
  2. (ii)
    for any $\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$, the Pythagorean theorem holds in the sense of
    $\langle\mathcal{T}f,\mathcal{T}f\rangle_{N}+\langle\mathcal{T}f-f,\mathcal{T}f-f\rangle_{N}=\langle f,f\rangle_{N}.$
Example 3.1
Hard thresholding operator $\mathcal{H}_{n}^{\lambda}$ and hyperinterpolation operator $\mathcal{L}_{n}$ belong to $\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$. Indeed, it is easy to check that all $\mathcal{H}_{n}^{\lambda}$ and $\mathcal{L}_{n}$ form an algebra. They satisfy the Pythagorean theorem which can be found in [2, Lemma 3.1] and [28, Lemma 5], respectively.
Proposition 3.2
If $\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$, then the following holds
  1. (i)
    $\langle\mathcal{T}f,f\rangle_{N}=\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}^{2}$;
  2. (ii)
    $\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}\leq\|f\|_{\ell^{2}(\mathbf{w})}$.
Proof.  (i) Since $\mathcal{T}$ satisfies Pythagorean theorem, it is clear that
$2\langle\mathcal{T}f,\mathcal{T}f\rangle_{N}-2\langle\mathcal{T}f,f\rangle_{N}+\langle f,f\rangle_{N}=\langle f,f\rangle_{N},$
which implies that
$\langle\mathcal{T}f,f\rangle_{N}=\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}^{2}.$
(ii) From (i), we have
$\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}^{2}=\langle\mathcal{T}f,f\rangle_{N}\leq\|f\|_{\ell^{2}(\mathbf{w})}\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})},$
where the inequality follows from the Cauchy-Schwarz inequality. $\square$
Lemma 3.2
If $\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ and $\mathcal{T}^{2}=\mathcal{T}$, then $\|\mathcal{T}\|_{\ell^{2}(\mathbf{w})}=1$.
Proof.  By direct computation, we have
$\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{T}^{2}f\|_{\ell^{2}(\mathbf{w})}=\left\|\mathcal{T}\left(\frac{\mathcal{T}f}{\left\|\mathcal{T}f\right\|_{\ell^{2}(\mathbf{w})}}\right)\right\|_{\ell^{2}(\mathbf{w})}\left\|\mathcal{T}f\right\|_{\ell^{2}(\mathbf{w})}\leq\|\mathcal{T}\|_{\ell^{2}(\mathbf{w})}\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}.$
Then combining the result (ii) in Proposition 3.2, we deduce that $\|\mathcal{T}\|_{\ell^{2}(\mathbf{w})}=1$. $\square$
Theorem 3.4
If $\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$ and if $\langle\mathcal{T}f,f\rangle_{N}=0$ for every $f\in\mathcal{C}(\mathbb{S}^{d})$, then $\mathcal{T}=0$.
Proof.  From Proposition 3.2 (i), we have
$\langle\mathcal{T}f,f\rangle_{N}=\|\mathcal{T}f\|_{\ell^{2}(\mathbf{w})}^{2}=0$
for all $f$. Thus, we deduce that $\mathcal{T}f=0$. $\square$
Corollary 3.1
If $\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$, $\mathcal{S}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$,and
$\langle\mathcal{T}f,f\rangle_{N}=\langle\mathcal{S}f,f\rangle_{N}$
for every $f\in\mathcal{C}(\mathbb{S}^{d})$, then $\mathcal{S}=\mathcal{T}$.
3.3 The product of hyper projection operators
In a Hilbert space, we often decompose the normal operators into several projection operators, which is very important in the spectral theorem [27]. Thus, we will inevitably encounter the operations of projection operators. Similarly, we first discuss the product of hyper projection operators:
Theorem 3.5
For two hyper projection operators $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the product $\mathcal{T}_{1}\mathcal{T}_{2}$ is also a hyper projection operators if and only if $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are commutative.
Proof.  Sufficiency. If $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are commutative, then for any $f,g\in\mathcal{C}(\mathbb{S}^{d})$
$\langle\mathcal{T}_{1}\mathcal{T}_{2}f,g\rangle_{N}=\langle\mathcal{T}_{2}f,\mathcal{T}_{1}g\rangle_{N}=\langle f,\mathcal{T}_{2}\mathcal{T}_{1}g\rangle_{N}=\langle f,\mathcal{T}_{1}\mathcal{T}_{2}g\rangle_{N},$
which implies that $\mathcal{T}_{1}\mathcal{T}_{2}$ is hyper self-adjoint. It is clear that
$(\mathcal{T}_{1}\mathcal{T}_{2})^{2}=\mathcal{T}_{1}\mathcal{T}_{2}\mathcal{T}_{1}\mathcal{T}_{2}=(\mathcal{T}_{1})^{2}(\mathcal{T}_{2})^{2}=\mathcal{T}_{1}\mathcal{T}_{2}.$
Necessity. Since $\mathcal{T}_{1}\mathcal{T}_{2}$ is a hyper projection operator, then $\mathcal{T}_{1}\mathcal{T}_{2}=(\mathcal{T}_{1}\mathcal{T}_{2})^{\ast}$ and
$\langle\mathcal{T}_{1}\mathcal{T}_{2}f,g\rangle_{N}=\langle\mathcal{T}_{2}f,\mathcal{T}_{1}g\rangle_{N}=\langle f,\mathcal{T}_{2}\mathcal{T}_{1}g\rangle_{N}=\langle f,\mathcal{T}_{1}\mathcal{T}_{2}g\rangle_{N}.$
Therefore, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are commutative. $\square$
Example 3.2
For two hyperinterpolation operators $\mathcal{L}_{n}$ and $\mathcal{L}_{m}$ with $n\leq m$, we have $\mathcal{L}_{n}\mathcal{L}_{m}=\mathcal{L}_{m}\mathcal{L}_{n}=\mathcal{L}_{n}$.
3.4 The sum of hyper projection operators
Theorem 3.6
For two hyper projection operators $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the sum $\mathcal{T}_{1}+\mathcal{T}_{2}$ is also a hyper projection operator if and only if $\mathcal{T}_{1}\mathcal{T}_{2}=0$ (or $\mathcal{T}_{2}\mathcal{T}_{1}=0)$
Proof.  Necessity. If $\mathcal{T}_{1}+\mathcal{T}_{2}$ is a hyper projection operator, then
$\mathcal{T}_{1}+\mathcal{T}_{2}$ $=(\mathcal{T}_{1}+\mathcal{T}_{2})^{2}=(\mathcal{T}_{1})^{2}+\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{2}\mathcal{T}_{1}+(\mathcal{T}_{2})^{2},$
$=\mathcal{T}_{1}+\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{2}\mathcal{T}_{1}+\mathcal{T}_{2},$
which implies that $\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{2}\mathcal{T}_{1}=0$. By direct computation, we have
$\left\{\begin{array}[]{c}\mathcal{T}_{1}\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{1}\mathcal{T}_{2}\mathcal{T}_{1}=0,\\ \mathcal{T}_{1}\mathcal{T}_{2}\mathcal{T}_{1}+\mathcal{T}_{2}\mathcal{T}_{1}\mathcal{T}_{1}=0,\end{array}\right.$
which means that
$\mathcal{T}_{1}\mathcal{T}_{2}=-\mathcal{T}_{1}\mathcal{T}_{2}\mathcal{T}_{1}=\mathcal{T}_{2}\mathcal{T}_{1}.$
Thus, $\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{2}\mathcal{T}_{1}=0$.
Sufficiency. If $\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{2}\mathcal{T}_{1}=0$, then
$\mathcal{T}_{1}+\mathcal{T}_{2}=(\mathcal{T}_{1}+\mathcal{T}_{2})^{2}=(\mathcal{T}_{1})^{2}+\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{2}\mathcal{T}_{1}+(\mathcal{T}_{2})^{2}=\mathcal{T}_{1}+\mathcal{T}_{2}.$
Thus, we have completed the proof. $\square$
Example 3.3
For any $f\in\mathcal{C}(\mathbb{S}^{d})$, consider the following two hyper projection operators
$\mathcal{L}_{n}~{}f=\sum_{\ell=0}^{n}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y^{(d)}_{\ell k}\right\rangle_{N}Y^{(d)}_{\ell k}\quad\text{and}\quad\mathcal{L}_{m-n}^{{}^{\prime}}~{}f=\sum_{\ell=n+1}^{m}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y^{(d)}_{\ell k}\right\rangle_{N}Y^{(d)}_{\ell k},\quad(n<m),$
then we have
$\mathcal{L}_{n}+\mathcal{L}_{m-n}^{{}^{\prime}}=\sum_{\ell=0}^{m}\sum_{k=1}^{Z(d,\ell)}\left\langle f,Y^{(d)}_{\ell k}\right\rangle_{N}Y^{(d)}_{\ell k}.$
3.5 The difference of hyper projection operators
Definition 3.8
For two hyper projection operators $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we call $\mathcal{T}_{1}$ is a suboperator of $\mathcal{T}_{2}$, if $\mathcal{T}_{1}\mathcal{C}(\mathbb{S}^{d})\subset\mathcal{T}_{2}\mathcal{C}(\mathbb{S}^{d})$.
Lemma 3.3
For two hyper projection operators $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, $\mathcal{T}_{1}$ is a suboperator of $\mathcal{T}_{2}$ if and only if one of the two conditions is satisfied:
  1. (i)
    $\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{2}\mathcal{T}_{1}=\mathcal{T}_{1}$;
  2. (ii)
    for any $f\in\mathcal{C}(\mathbb{S}^{d})$, $\|\mathcal{T}_{1}f\|_{\ell^{2}(\mathbf{w})}\leq\|\mathcal{T}_{2}f\|_{\ell^{2}(\mathbf{w})}$
Proof.  Necessity. For any $f\in\mathcal{C}(\mathbb{S}^{d})$, since $\mathcal{T}_{1}f\in M_{1}$, where $M_{1}$ is the image of $\mathcal{T}_{1}$, we have $\mathcal{T}_{2}\mathcal{T}_{1}f=\mathcal{T}_{1}f$. Thus, we obtain
$\mathcal{T}_{2}\mathcal{T}_{1}=\mathcal{T}_{1}.$
Furthermore, it is obvious that
$\langle\mathcal{T}_{2}\mathcal{T}_{1}f,g\rangle_{N}=\langle f,\mathcal{T}_{1}\mathcal{T}_{2}g\rangle_{N}=\langle\mathcal{T}_{1}f,g\rangle_{N}=\langle f,\mathcal{T}_{1}g\rangle_{N},$
which implies that $\mathcal{T}_{2}\mathcal{T}_{1}=\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{1}$. By direct computation and Lemma 3.2, we have
$\|\mathcal{T}_{1}f\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{T}_{1}\mathcal{T}_{2}f\|_{\ell^{2}(\mathbf{w})}\leq\|\mathcal{T}_{1}\|_{\ell^{2}(\mathbf{w})}\|\mathcal{T}_{2}f\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{T}_{2}f\|_{\ell^{2}(\mathbf{w})}.$
Sufficiency. If $\mathcal{T}_{1}$ is not a suboperator of $\mathcal{T}_{2}$, then there exist an element $\psi_{0}\in M_{1}$ but $\psi_{0}\notin M_{2}$, where $M_{1},M_{2}$ are the images of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, respectively. Let $\hat{\psi}_{0}$ be the orthogonal projection of $\psi_{0}$ in $M_{2}$. Then $\|\hat{\psi}_{0}\|_{\ell^{2}(\mathbf{w})}<\|{\psi}_{0}\|_{\ell^{2}(\mathbf{w})}$ and $\mathcal{T}_{2}\psi_{0}=\hat{\psi}_{0}$. Thus,
$\|\mathcal{T}_{2}\psi_{0}\|_{\ell^{2}(\mathbf{w})}=\|\hat{\psi}_{0}\|_{\ell^{2}(\mathbf{w})}<\|{\psi}_{0}\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{T}_{1}\psi_{0}\|_{\ell^{2}(\mathbf{w})},$
which contradicts the assumption. Therefore, $\mathcal{T}_{1}$ is a suboperator of $\mathcal{T}_{2}$. $\square$
Remark 3.6
Since filtered hyperinterpolation operator $\mathcal{F}_{n,N}$, hard thresholding hyperinterpolation operator $\mathcal{H}_{n}^{\lambda}$ and Lasso hyperinterpolation operator $\mathcal{L}_{n}^{\lambda}$ are not hyper projection operators, they fail to be suboperators of hyperinterpolation operators even if satisfing the conditions (i) or (ii) in Lemma 3.3.
Next, we focus on the difference between two hyper projection operators:
Theorem 3.7
For two hyper projection operators $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the difference $\mathcal{T}_{2}-\mathcal{T}_{1}$ is a hyper projection operator if and only if $\mathcal{T}_{1}$ is a suboperator of $\mathcal{T}_{2}$.
Proof.  Necessity. If $\mathcal{T}_{2}-\mathcal{T}_{1}$ is a hyper projection operator, then we have
$\langle\mathcal{T}_{2}(\mathcal{T}_{2}-\mathcal{T}_{1})f,g\rangle_{N}=\langle(\mathcal{T}_{2}-\mathcal{T}_{2}\mathcal{T}_{1})f,g\rangle_{N}=\langle f,(\mathcal{T}_{2}-\mathcal{T}_{1})\mathcal{T}_{2}g\rangle_{N}=\langle f,(\mathcal{T}_{2}-\mathcal{T}_{1}\mathcal{T}_{2})g\rangle_{N}.$
Thus we can deduce that $\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{2}\mathcal{T}_{1}$. Besides, the following holds:
$(\mathcal{T}_{2}-\mathcal{T}_{1})^{2}=\mathcal{T}_{2}^{2}-\mathcal{T}_{2}\mathcal{T}_{1}-\mathcal{T}_{1}\mathcal{T}_{2}+\mathcal{T}_{1}^{2}=\mathcal{T}_{2}-\mathcal{T}_{1},$ (3.8)
which implies that $\mathcal{T}_{2}\mathcal{T}_{1}=\mathcal{T}_{1}\mathcal{T}_{2}=\mathcal{T}_{1}$.
Sufficiency. By Lemma 3.3 (i), the equation (3.8) holds. On the flip side, we have
$\langle(\mathcal{T}_{2}-\mathcal{T}_{1})f,g\rangle_{N}=\langle\mathcal{T}_{2}f,g\rangle_{N}-\langle\mathcal{T}_{1}f,g\rangle_{N}=\langle f,\mathcal{T}_{2}g\rangle_{N}-\langle f,\mathcal{T}_{1}g\rangle_{N}=\langle f,(\mathcal{T}_{2}-\mathcal{T}_{1})g\rangle_{N}.$
Thus, we have completed the proof. $\square$
Example 3.4
For two hyperinterpolation operators $\mathcal{L}_{n}$ and $\mathcal{L}_{m}$ with $n\leq m$, $\mathcal{L}_{n}$ is a suboperator of $\mathcal{L}_{m}$. Thus, the difference $\mathcal{L}_{m}-\mathcal{L}_{n}$ is also a hyper projection operator. Specifically, we have
$(\mathcal{L}_{m}-\mathcal{L}_{n})f=\sum_{\ell=n+1}^{m}\sum_{k=1}^{Z(d,\ell)}\left\langle{f,Y_{\ell k}^{(d)}}\right\rangle_{N}Y_{\ell k}^{(d)}.$
4 Ideals and involutions
In this section, we mainly investigate the ideals and involutions of hyperinterpolation operators. From the above, we know that both hard thresholding hyperinterpolation operator $\mathcal{H}_{n}^{\lambda}$ and hyperinterpolation operator $\mathcal{L}_{n}$ are elements of the hyper algebra $\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))$. Now we define three sets that can be useful in the latter,
$\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d})):=\{\mathcal{T}:\mathcal{T}=\mathcal{L}_{n}\text{ or }\mathcal{T}=\mathcal{H}_{n}^{\lambda},\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))\}$
and
$\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d})):=\{\mathcal{T}:\mathcal{T}=\mathcal{H}_{n}^{\lambda},\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))\},$
and
$\mathfrak{L}(\mathcal{C}(\mathbb{S}^{d})):=\{\mathcal{T}:\mathcal{T}=\mathcal{L}_{n},\mathcal{T}\in\mathfrak{B}(\mathcal{C}(\mathbb{S}^{d}))\}.$
4.1 Ideals
We review the definition of an ideal:
Definition 4.1
[27, Definition 11.1, p.275] A subset $\mathcal{S}$ of a commutative complex algebra $\mathcal{A}$ is said to be an ideal if
  1. (i)
    $\mathcal{S}$ is a subspace of $\mathcal{A}$;
  2. (ii)
    $xy\in\mathcal{S}$ whenever $x\in\mathcal{A}$ and $y\in\mathcal{S}$.
If $\mathcal{S}\neq\mathcal{A}$, $\mathcal{S}$ is a proper ideal. Maximal ideals are proper ideals that are not contained in any larger proper ideal [27]. Then we have the following important result.
Theorem 4.1
The subset $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is a maximal ideal of $\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$.
Proof.  It is clear that $\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$ and $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ are commutative hyper algebra, respectively. And we can easily verify that $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is a subspace of $\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$.
We first prove that $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is a proper ideal. Let $\mathcal{H}_{n}^{\lambda}\in\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ and $\mathcal{H}_{m}^{\lambda},\mathcal{L}_{m}\in\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$, where $m$ may be not equal to $n$. Then we have
$\mathcal{L}_{m}\mathcal{H}_{n}^{\lambda}=\left\{\begin{array}[]{ll}\mathcal{H}_{n}^{\lambda},&m\geq n,\\ \mathcal{H}_{m}^{\lambda},&m<n,\end{array}\right.$
and
$\mathcal{H}_{m}^{\lambda}\mathcal{H}_{n}^{\lambda}=\left\{\begin{array}[]{ll}\mathcal{H}_{n}^{\lambda},&m\geq n,\\ \mathcal{H}_{m}^{\lambda},&m<n.\end{array}\right.$
Indeed, for $\mathcal{L}_{m}\mathcal{H}_{n}^{\lambda}$ and $f\in\mathcal{C}(\mathbb{S}^{d})$, we obtain
$\mathcal{L}_{m}\mathcal{H}_{n}^{\lambda}f$ $=\sum_{\ell=0}^{m}\sum_{k=1}^{Z(d,\ell)}\left\langle\sum_{\ell^{\prime}=0}^{n}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\eta_{H}\left(\left\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\right\rangle_{N},\lambda\right)Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}$
$=\sum_{\ell=0}^{m}\sum_{k=1}^{Z(d,\ell)}\sum_{\ell^{\prime}=0}^{n}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\eta_{H}\left(\left\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\right\rangle_{N},\lambda\right)\left\langle Y_{\ell^{\prime}k^{\prime}}^{(d)},Y_{\ell k}^{(d)}\right\rangle_{N}Y_{\ell k}^{(d)}.$
Similarly, one can prove that
$\mathcal{H}_{m}^{\lambda}\mathcal{H}_{n}^{\lambda}f=\sum_{\ell=0}^{m}\sum_{k=1}^{Z(d,\ell)}\eta_{H}\left(\sum_{\ell^{\prime}=0}^{n}\sum_{k^{\prime}=1}^{Z(d,\ell^{\prime})}\eta_{H}\left(\left\langle f,Y_{\ell^{\prime}k^{\prime}}^{(d)}\right\rangle_{N},\lambda\right)\left\langle Y_{\ell k}^{(d)},Y_{\ell^{\prime}k^{\prime}}^{(d)}\right\rangle_{N},\lambda\right)Y_{\ell k}^{(d)}.$
Now, we prove that $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is maximal. If $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is contained in a larger set $\mathfrak{H}^{\prime}(\mathcal{C}(\mathbb{S}^{d}))\subsetneqq\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$. Then we can deduce that $\mathfrak{H}^{\prime}(\mathcal{C}(\mathbb{S}^{d}))$ contains some hyperinterpolation operators $\mathcal{L}_{n}$. Then there exists a hyperinterpolation operator $\mathcal{L}_{k}\in\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$ but $\mathcal{L}_{k}\notin\mathfrak{H}^{\prime}(\mathcal{C}(\mathbb{S}^{d}))$ such that
$\mathcal{L}_{k}\mathcal{L}_{n}=\mathcal{L}_{k}\notin\mathfrak{H}^{\prime}(\mathcal{C}(\mathbb{S}^{d})).$
Thus, we have completed the proof. $\square$
4.2 Involutions
We first give the definition of involution:
Definition 4.2
[27, Definition 11.4, p.287] A mapping $x\to x^{\ast}$ of a complex (not necessarily commutative) algebra $\mathcal{A}$ into $\mathcal{A}$ is called an involution on $\mathcal{A}$ if it has following four properties, for all $x\in\mathcal{A}$, $y\in\mathcal{A}$, and $\alpha\in\mathbb{C}$:
  1. (i)
    $(x+y)^{\ast}=x^{\ast}+y^{\ast}$;
  2. (ii)
    $(\alpha x)^{\ast}=\bar{\alpha}x^{\ast}$;
  3. (iii)
    $(xy)^{\ast}=y^{\ast}x^{\ast}$;
  4. (iv)
    $x^{\ast\ast}=x$.
If $x^{\ast}=x$ for any $x\in\mathcal{A}$, then $x$ is called hermitian, or self-adjoint [27]. Then we can define the hyper $C^{\ast}$-algebra:
Definition 4.3
A hyper algebra $\mathcal{A}$ with an involution $x\to x^{\ast}$ that satisfies
$\|xx^{\ast}\|_{\ell^{2}(\mathbf{w})}=\|x\|_{\ell^{2}(\mathbf{w})}^{2}$
for every $x\in\mathcal{A}$ is called a hyper $C^{\ast}$-algebra.
Proposition 4.1
The hyper algebra $\mathfrak{L}(\mathcal{C}(\mathbb{S}^{d}))$ is a hyper $C^{\ast}$-algebra.
Proof.  Let $\mathcal{L}_{n}\in\mathfrak{L}(\mathcal{C}(\mathbb{S}^{d}))$. Since $\mathcal{L}_{n}$ is a hyper projection operator, it is hyper self-adjoint. It is obvious that $\mathcal{L}_{n}=(\mathcal{L}_{n})^{\ast}$ which means $\mathcal{L}_{n}(\mathcal{L}_{n})^{\ast}=\mathcal{L}_{n}$. By Lemma 3.2, we have
$\|\mathcal{L}_{n}(\mathcal{L}_{n})^{\ast}\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{L}_{n}\|_{\ell^{2}(\mathbf{w})}=\|\mathcal{L}_{n}\|_{\ell^{2}(\mathbf{w})}^{2}=1.$
Thus, we have completed the proof. $\square$
Remark 4.1
For every $\mathcal{H}_{n}^{\lambda}\in\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$, since $\mathcal{H}_{n}^{\lambda}$ is not hyper adjoint , the hyper algebra $\mathfrak{H}(\mathcal{C}(\mathbb{S}^{d}))$ is not a hyper $C^{\ast}$-algebra.
Then we give the definition of the hyper homomorphism:
Definition 4.4
Suppose $\mathcal{A}$ is a hyper algebra and $\phi$ is a linear operator on $\mathcal{A}$ which is not identically zero. If
$\phi(xy)=\phi(x)\phi(y)$
for all $x,y\in\mathcal{A}$, then $\phi$ is called a hyper homomorphism on $\mathcal{A}$.
Theorem 4.2
Hyperinterpolation operator $\mathcal{L}_{n}$ is a hyper homomorphism on the hyper algebra $\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$.
Proof.  Let $\mathcal{H}_{m_{1}}^{\lambda},\mathcal{H}_{m_{2}}^{\lambda},\mathcal{L}_{s_{1}},\mathcal{L}_{s_{2}}\in\mathfrak{I}(\mathcal{C}(\mathbb{S}^{d}))$. Without loss of generality, we assume $m_{i},s_{i}\leq n$ for $i=1,2$. Elementary calculations yield the following three cases
  1. (i)
    for $\mathcal{L}_{s_{1}},\mathcal{L}_{s_{2}}$, we have
    $\mathcal{L}_{n}(\mathcal{L}_{s_{1}}\mathcal{L}_{s_{2}})=\left\{\begin{array}[]{ll}\mathcal{L}_{n}(\mathcal{L}_{s_{1}})=\mathcal{L}_{s_{1}}=\mathcal{L}_{n}(\mathcal{L}_{s_{1}})\mathcal{L}_{n}(\mathcal{L}_{s_{2}}),&s_{1}<s_{2},\\ \mathcal{L}_{n}(\mathcal{L}_{s_{2}})=\mathcal{L}_{s_{2}}=\mathcal{L}_{n}(\mathcal{L}_{s_{1}})\mathcal{L}_{n}(\mathcal{L}_{s_{2}}),&s_{1}\geq s_{2}.\end{array}\right.$
  2. (ii)
    for $\mathcal{L}_{s_{1}},\mathcal{H}_{m_{1}}^{\lambda}$, we have
    $\mathcal{L}_{n}(\mathcal{L}_{s_{1}}\mathcal{H}_{m_{1}}^{\lambda})=\left\{\begin{array}[]{ll}\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda})=\mathcal{H}_{m_{1}}^{\lambda}=\mathcal{L}_{n}(\mathcal{L}_{s_{1}})\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda}),&m_{1}<s_{1},\\ \mathcal{L}_{n}(\mathcal{H}_{s_{1}}^{\lambda})=\mathcal{H}_{s_{1}}^{\lambda}=\mathcal{L}_{n}(\mathcal{L}_{s_{1}})\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda}),&m_{1}\geq s_{1}.\end{array}\right.$
    The above relation for $\mathcal{H}_{m_{1}}^{\lambda}\mathcal{L}_{s_{1}}$ also holds since hard thresholding operator and hyperinterpolation operator are commutative.
  3. (iii)
    for $\mathcal{H}_{m_{1}}^{\lambda},\mathcal{H}_{m_{2}}^{\lambda}$, we have
    $\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda}\mathcal{H}_{m_{2}}^{\lambda})=\left\{\begin{array}[]{ll}\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda})=\mathcal{H}_{m_{1}}^{\lambda}=\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda})\mathcal{L}_{n}(\mathcal{H}_{m_{2}}^{\lambda}),&m_{1}<m_{2},\\ \mathcal{L}_{n}(\mathcal{H}_{m_{2}}^{\lambda})=\mathcal{H}_{m_{2}}^{\lambda}=\mathcal{L}_{n}(\mathcal{H}_{m_{1}}^{\lambda})\mathcal{L}_{n}(\mathcal{H}_{m_{2}}^{\lambda}),&m_{1}\geq m_{2}.\end{array}\right.$
Thus, we have completed the proof. $\square$
Definition 4.5
An operator $\mathcal{T}$ is called a hyper positive operator if
$\langle\mathcal{T}f,f\rangle_{N}\geq 0$
for every $f\in\mathcal{C}(\mathbb{S}^{d})$.
Example 4.1
Hyperinterpolation operator $\mathcal{L}_{n}$, hard thresholding hyperinterpolation operator $\mathcal{H}_{n}^{\lambda}$, Lasso hyperinterpolation operator $\mathcal{L}_{n}^{\lambda}$ and filtered hyperinterpolation operator $\mathcal{F}_{n,N}$ are hyper positive operators. However, the generalized hyperinterpolation operator $\mathfrak{G}\mathfrak{L}_{n}$ may not be a positive operator. Indeed, for any $f\in\mathcal{C}(\mathbb{S}^{d})$, we have
$\langle\mathfrak{G}\mathfrak{L}_{n}f,f\rangle_{N}=\left\langle\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}Y_{\ell k}^{(d)},f\right\rangle_{N}=\sum_{\ell=0}^{n}a_{n\ell}\sum_{k=1}^{Z(d,\ell)}\langle f,Y_{\ell k}^{(d)}\rangle_{N}^{2}.$
We can not judge the sign of $\langle\mathfrak{G}\mathfrak{L}_{n}f,f\rangle_{N}$ due to the existence of $a_{n\ell}$’s.
5 Final remark
In this paper, our focus has been on investigating the algebraic properties of hyperinterpolation-class on the unit sphere, particularly emphasizing hyper projection operators, ideals, and involutions properties, under classical quadrature exactness assumptions (2.7). Our main contributions can be summarized in three key aspects:
  1. $\bullet$
    Establishment of hyper projection operators: We developed corresponding theories of hyper projection operators in the context of a semi (inner) product. This provides a solid foundation for understanding the projection properties of hyperinterpolation-class.
  2. $\bullet$
    Introduction of hyper algebra: We introduced the concept of a hyper algebra and demonstrated that the hyper algebra containing hard thresholding hyperinterpolation operators forms the maximal ideal within the broader hyper algebra comprised of both hard thresholding hyperinterpolation and hyperinterpolation operators. This elucidates the algebraic structures underlying hyperinterpolation-class.
  3. $\bullet$
    Role of involutions in hyper algebra: We highlighted the significance of involutions in hyper algebra, revealing their connection to hyper $C^{\ast}$-algebras and hyper homomorphism. This sheds light on the structural properties of hyperinterpolation-class within an algebraic framework.
Looking ahead, leveraging the Marcinkiewicz-Zygmund inequality [11, 17, 22, 36] offers a promising avenue for further exploration of the algebraic properties of hyperinterpolation-class, particularly in scenarios where quadrature exactness assumptions may not hold [6]. By relaxing these assumptions, we can extend our understanding and applicability of hyperinterpolation.
Furthermore, given that we have established the concepts of hyper $C^{\ast}$-algebras and hyper homomorphism, future research could delve into properties of normal operators and spectral theorems within the framework of hyperinterpolation. This opens up opportunities to explore spectral properties and analyze the behavior of operators in hyperinterpolation settings, potentially leading to deeper insights and applications in various fields.